OpenStudy (jbolyefo):

Please help, I'll fan and medal Let u = <6, -9>, v = <1, -4>. Find u - v.

OpenStudy (jbolyefo):

@Sila1453 can you help?

OpenStudy (jbolyefo):

@pooja195

OpenStudy (jbolyefo):

@IMStuck @mathmate

OpenStudy (mww):

Vector subtraction is quite intuitive. Just align elements and subtract! i.e. \[\left(\begin{matrix}a \\ b\end{matrix}\right) - \left(\begin{matrix}c \\ d\end{matrix}\right)=\left(\begin{matrix}a-c \\ b-d\end{matrix}\right)\]

OpenStudy (jbolyefo):

I think I did it wrong, I got -5/12 @mww

OpenStudy (jbolyefo):

I don't know what numbers to put where

OpenStudy (welshfella):